OpenStudy (anonymous):

2.3 is 46% of what number?

OpenStudy (abb0t):

This should be posted in the mathematics section, not chemistry. But, either way, you should know for chemistry as this is important also. Percents are essentially ratios, because they are out of 100, usually. So you have 46% = \(\sf \frac{46}{100}\) = 0.46 Now, you could either set up a proportion, OR simply multiply. 2.3 \(\times\) 0.46 = _______ that is how much 46% of 2.3 is.

OpenStudy (cuanchi):

I think the question is in other direction it is asking what is the value of the 100% if the 46% is equal to 2.3 %= Part x 100/Total => Total = Part x 100/% Total= 2.3 x 100/46%= 5